OpenStudy (anonymous):

The measure of each exterior angle of a regular polygon is 40°. How many sides does the polygon have?

OpenStudy (anonymous):

@jim_thompson5910 IS THERE AN EQUATION TO SOLVE THIS?

OpenStudy (anonymous):

sorry for the caps =^.^=

jimthompson5910 (jim_thompson5910):

let n = # of sides E = measure of exterior angle

jimthompson5910 (jim_thompson5910):

the formula is E = 360/n solving for n gives you n*E = 360 n = 360/E

OpenStudy (anonymous):

9 :)

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

AWESOME!

jimthompson5910 (jim_thompson5910):

lol love the enthusiasm and I'm glad it's all clicking now